# FC-Attack: Jailbreaking Multimodal Large Language Models via Auto-Generated Flowcharts

### v2

@@ Introduction
With the advancement of Large Language Models (LLMs), Multimodal Large Language Models (MLLMs) that integrate vision (images and videos) and text, such as GPT-4o [18] and Qwen2.5-VL [6], have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding [24, 20].
@@
Despite being powerful, recent studies [14, 32] have revealed that MLLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards and gain harmful knowledge. Such vulnerabilities pose remarkable safety risks to the Internet and the physical world. For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion [36]. To better safeguard MLLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al. [52] providing a quantitative understanding regarding the adversarial vulnerability of MLLMs. Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
@@
Currently, jailbreak attacks against MLLMs can be broadly categorized into two main types: optimization-based attacks [8, 23] and prompt-based attacks [14, 39]. Optimization-based attacks use white-box gradient methods to craft adversarial perturbations on visual prompt aligned with harmful text. They are effective but slow and have limited transferability in black-box scenarios. In contrast, prompt-based jailbreaks require only black-box access and work by injecting malicious visual cues into benign prompts to exploit MLLMs’ text-focused safety alignment.
@@
To better improve the attack transferability and its effectiveness, we propose a novel prompt-based jailbreak attack, namely $\mathsf{FC\text{-}Attack}$. Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts (images and videos) as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards. Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and fine-tune a pre-trained LLM to obtain a step-description generator. (2) Jailbreak Deployment: This stage uses the generator to produce steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts. Together with the benign textual prompt, the visual prompt is fed into MLLMs to achieve the jailbreak. Note that the harmful flowcharts are generated automatically without hand-crafted effort.
@@
Our evaluation on the Advbench dataset shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5. Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways. To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of MLLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure 1), numbers of descriptions, and font styles in flowcharts. These experiments show that MLLMs exhibit higher safety in the text modality but weaker in the visual modality. Moreover, we find that even flowcharts with a one-step harmful description can achieve high ASR, as evidenced by the Gemini-1.5 model, where the ASR reaches $86\%$. Furthermore, font styles in flowcharts also contribute to the ASR increase. For instance, when the font style is changed from “Times New Roman” to “Pacifico”, the ASR increases from $4\%$ to $28\%$ on the model with the lowest ASR (Claude-3.5) under the original style. To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision [28], JailGuard [47], AdaShield-S [38], and AdaShield-A [38]. Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$. However, it also reduces MLLM’s utility on benign datasets, which calls for more effective defenses.
@@
In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities [51, 10], which have driven the emergence of MLLMs [46] (also known as Large Vision Language Models, LVLMs). MLLMs combine visual understanding with language comprehension, showing promising capabilities in visual downstream tasks, including Visual Question Answering (VQA) [3, 21, 33], image captioning [17, 22], and visual commonsense reasoning [45, 35]. Notably, some MLLMs are capable of processing both image and video inputs, enabling broader applications across multimodal scenarios.
In this paper, we consider both popular open-source and widely used production MLLMs. These MLLMs are the most widely used, and all of them have been aligned to ensure safety. Detailed information are introduced in Appendix A.
@@
Similar to LLMs, which have been shown to be vulnerable to jailbreak attacks [42], MLLMs also remain susceptible despite safety alignment. Current attacks can be categorized into two types: optimization-based and prompt-based attacks. Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs. These methods typically require white-box access to the model, where they obtain the output logits of MLLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts [4, 34, 31] (e.g., the target can be “Sure! I’m ready to answer your question.”). Carlini et al. [9] are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs. Building on this, Bailey et al. [8] introduce the Behaviour Matching Algorithm, which trains adversarial images to make MLLMs output behavior that matches a target in specific contextual inputs. This process requires the model’s output logits to align closely with those of the target behavior. Additionally, they propose Prompt Matching, where images are used to induce the model to respond to specific prompts. Li et al. [23] take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking. Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4. This approach enhances the harmfulness of the images, enabling more effective attacks.
@@
Unlike optimization-based attacks, prompt-based attacks only need black-box access to successfully attack the model without introducing adversarial perturbations into images. Gong et al. [14] discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities. To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content. This method can effectively attack MLLMs without requiring any training. Wang et al. [39] identifies a phenomenon named Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs. Specifically, even if harmful instructions in text or images are rearranged, MLLMs can still correctly interpret their meaning. However, the safety mechanisms of MLLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs. Compared to optimization-based attacks, prompt-based attacks usually achieve higher success rates against closed-source models. Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access.
@@
Adversary’s Goal. The adversary’s goal is to exploit attacks to bypass the protective mechanisms of MLLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy [30]. This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of MLLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort. These objectives pose severe societal impacts and risks to the model providers.
@@
In this section, we introduce the framework of $\mathsf{FC\text{-}Attack}$ (as shown in Figure 2), which consists of two stages: Step-Description Generator Building and Jailbreak Deployment.
@@
After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description. This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz [16]). Together with a benign textual prompt $p_{b}$ (more details are in Appendix B), the flowchart will be fed into the aligned MLLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
@@
Target Model. We test $\mathsf{FC\text{-}Attack}$ on seven popular MLLMs, including the open-source models Llava-Next ( llama3-llava-next-8b) [26], Qwen2-VL (Qwen2-VL-7B-Instruct) [37], and InternVL-2.5 (InternVL-2.5-8B) [12] as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) [29], GPT-4o (gpt-4o-2024-08-06) [18], Claude (claude-3-5-sonnet-20240620) [2], and Gemini (gemini-1.5-flash) [15]. Moreover, we also test $\mathsf{FC\text{-}Attack}$ on LLMs via video, including Qwen-vl-max (Qwen-vl-max-latest) [7],Qwen2.5-Omni [40] and LLaVA-Video-7B-Qwen2 [49].
Dataset. Following Chao et al. [11], we utilize the deduplicated version of AdvBench [54], which includes 50 representative harmful queries. Based on AdvBench, we use $\mathsf{FC\text{-}Attack}$ to generate $3$ types of flowcharts for each harmful query, which includes $150$ jailbreak flowcharts in total. To assess whether defense methods have the critical issue of “over-defensiveness” when applied to benign datasets, we utilize a popular evaluation benchmark, MM-Vet [43].
@@
HADES $4$ $16$ $2$ $2$ $20$ $10$ $8$
SI-Attack $36$ $14$ 00 $69$ $24$ $42$ $40$
MM-SafetyBench 00 00 00 $50$ $50$ $54$ $16$
VA-Jailbreak $6$ $18$ $2$ $2$ $40$ $22$ $16$
FigStep 00 $2$ 00 $30$ $62$ $36$ 00
Ours (Vertical) $8$ $8$ 00 $76$ $76$ $84$ $68$
@@
ASR=# Queries Successfully Jailbroken# Original Harmful $\text{ASR}=\frac{\text{\# Queries Successfully Jailbroken}}{\text{\# Original Harmful Queries}}.$ (5)
Following the judge prompt [11], we employ GPT-4o to serve as the evaluator.
FC-Attack Deployment. Referring to Section 4, $\mathsf{FC\text{-}Attack}$ consists of two stages. For the Step-Description Generator Building, we first use GPT-4o to randomly generate several daily topics and $3$ few-shot examples, which are then combined into a prompt and fed into GPT-4o to construct the dataset $\mathcal{D}_{t}$. In our experiments, the number of descriptions in the flowchart is limited to a maximum of $10$ steps, as too many descriptions can result in excessive length in one direction of the image. The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix C). We then select Mistral-7B-Instruct-v0.1 [19] as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA. The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, $2$ epochs, a batch size of $8$, a learning rate of $51e-$ 5, and a weight decay of $51e-$ 5. For the jailbreak deployment stage, we set the temperature to $0.3$ for all MLLMs for a fair comparison.
Baselines. To validate the effectiveness of $\mathsf{FC\text{-}Attack}$, we adopt five jailbreak attacks as baselines, which are categorized into black-box attacks (MM-SafetyBench [27], SI-Attack [50], and FigStep [14]) and white-box attacks (HADES [23], VA-Jailbreak [31]).
For black-box attacks, MM-SafetyBench utilizes StableDiffusion [32] and GPT-4 [1] to generate harmful images and texts based on AdvBench. The input harmful images and texts used in SI-Attack are from the outputs of MM-SafetyBench, while FigStep is set up using their default settings [14].
For white-box attacks, all input data, including images and texts, is obtained from MM-SafetyBench’s outputs, with the attack step size uniformly set to 1/255. HADES employs LLaVa-1.5-7b [25] as the attack model, running 3,000 optimization iterations with a batch size of 2. For VA-Jailbreak, LLaVa-1.5-7b [25] is used as the attack model, setting the epsilon of the attack budget to 32/255, with 5,000 optimization iterations and a batch size of 8. To align with the black-box scenario considered in this paper, we adopt a model transfer strategy, where these white-box methods are trained on one model (LLava-1.5-7b) and then transferred to our target testing models.
@@
To mitigate the attacks, we explore several possible defense methods including Llama-Guard3-V, JailGuard, and AdaShield. Llama-Guard3-V (Llama-Guard-3-11B-Vision) [28] determines whether the input is safe by feeding both the image and text into the model. JailGuard [47] generates input variants and evaluates them using MiniGPT-4 [53], identifying harmful content by comparing differences in the responses. AdaShield-S employs static prompts in the textual prompt to defend against attacks, while AdaShield-A uses Vicunav1.5-13B as a defender to adaptively rewrite defensive prompts [38].
Figure 3: ASR under different prompts against MLLMs.
@@
In this section, we explore the performance of $\mathsf{FC\text{-}Attack}$ and conduct ablation study and defense research. We conduct jailbreak experiments on MLLMs for $\mathsf{FC\text{-}Attack}$. As shown in Figure A1 , it is a successful jailbreak case on Gemini-1.5.
@@
Creepster 14$\uparrow$ 24$\downarrow$ 8$\uparrow$ 94 90$\downarrow$ 90 90
Fruktur 18$\uparrow$ 28$\downarrow$ 18$\uparrow$ 98$\uparrow$ 86$\downarrow$ 90 88$\downarrow$
Pacifico 14$\uparrow$ 30 28$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 96$\uparrow$
Shojumaru 20$\uparrow$ 30 12$\uparrow$ 90$\downarrow$ 94$\uparrow$ 90 88$\downarrow$
UnifrakturMaguntia 12$\uparrow$ 24$\downarrow$ 26$\uparrow$ 90$\downarrow$ 90$\downarrow$ 90 92$\uparrow$
@@
Jailbreaking via Images. In Table 1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models. We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both models compared to all baselines. For example, the ASRs are $94\%$, $92\%$, $90\%$, and $90\%$ on Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, respectively. However, the ASR on some production models, such as Claude-3.5, GPT-4o, and GPT-4o mini, is relatively low, at $4\%$, $30\%$, and $10\%$, respectively. This might be because these production models have more advanced and updated visual safety alignment strategies.
@@
Jailbreaking via Videos. To conduct attacks from the video modality, we transform each jailbreak image into a 3-second video by setting all frames into the same image. Note that we also consider the Procedure Flowcharts, where each part (1 question and 5 steps) has been sequentially filled into a 0.5s video frame, resulting in a 3s video. We then evaluate the effectiveness of video jailbreak on three models: Qwen-vl-max, Qwen2.5-Omni and LLaVA-Video. The performance is summarized in Table 4. Our $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves a stable 88% ASR, whereas HADES peaks at 46% on Qwen-vl-max (dropping to 28% on LLaVA-Video) and Figstep fluctuates between 78% on Qwen-vl-max and 2% on Qwen2.5-Omni, highlighting our method’s consistent performance across models. As shown in Figure A2, jailbreaks using harmful text have an extremely low ASR. When the same harmful queries and steps are delivered via the video modality, the MLLMs become highly vulnerable, with ASR up to 88%.
@@
As shown in Figure 3 when using only the harmful query (text) as input, we observe very low ASR. The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively. This indicates that the textual modality of these MLLMs has relatively robust defenses against such inputs. However, when both the harmful query and the step descriptions are input as text, the ASR increases to $36\%$ on Gemini-1.5, and to $16\%$ and $6\%$ on Llava-Next and InternVL-2.5, respectively, while remaining at $0\%$ on the other models. When this information is converted into a flowchart and only a benign textual prompt is provided, the ASR on these models improves significantly. This demonstrates that the defenses of MLLMs in the visual modality have noticeable weaknesses compared with the language modality.
@@
Numbers of Steps in Flowcharts. As described in Section 4, flowcharts of $\mathsf{FC\text{-}Attack}$ are generated from step descriptions. In this section, we aim to explore the impact of the number of steps in flowcharts on jailbreak effectiveness. Therefore, we reduce the number of steps to 1, 3, and 5, respectively. Table 2 presents the ASR results for four types of flowcharts (Vertical, Horizontal, S-shaped, and Ensemble) with varying numbers of steps. We find that, even with only one step in the description, flowcharts achieve relatively high ASR. For example, for Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, the ASR for Ensemble at 1 step is $86\%$, $70\%$, $88\%$, and $82\%$, respectively. As the number of steps increases, the ASR for almost all flowchart types improves significantly. For instance, the Horizontal ASR of Gemini-1.5 increases from $78\%$ at “1 step” to $86\%$ at “3 steps” and $88\%$ at “5 steps”. Similarly, the S-shaped ASR of InternVL-2.5 improves from $68\%$ at “1 step” to $92\%$ at “5 steps”. This suggests that increasing the number of step descriptions makes the model more vulnerable and susceptible to jailbreak attacks.

### v3

@@ Introduction
With the advancement of Large Language Models (LLMs), Multimodal Large Language Models (MLLMs) that integrate vision (images and videos) and text, such as GPT-4o Hurst et al. (2024) and Qwen2.5-VL Bai et al. (2025a), have demonstrated emergent abilities and achieved impressive performance on downstream tasks related to visual understanding Liu et al. (2024a); Jin et al. (2024).
@@
Despite being powerful, recent studies Gong et al. (2025); Rombach et al. (2022) have revealed that MLLMs are vulnerable to jailbreak attacks whereby the adversary uses malicious methods to bypass safeguards and gain harmful knowledge. Such vulnerabilities pose remarkable safety risks to the Internet and the physical world. For instance, in January 2025, the world witnessed the first case where ChatGPT was used to conduct an explosion The Times (2025). To better safeguard MLLMs and proactively address their vulnerabilities, model researchers make many efforts in this regard, such as Zhao et al. (2023b) providing a quantitative understanding regarding the adversarial vulnerability of MLLMs. Previous studies often create adversarial datasets tailored to specific models, which tend to perform poorly on other models.
@@
Currently, jailbreak attacks against MLLMs can be broadly categorized into two main types: optimization-based attacks Bailey et al. (2024); Li et al. (2024b) and prompt-based attacks Gong et al. (2025); Wang et al. (2025). Optimization-based attacks use white-box gradient methods to craft adversarial perturbations on visual prompt aligned with harmful text. They are effective but slow and have limited transferability in black-box scenarios. In contrast, prompt-based jailbreaks require only black-box access and work by injecting malicious visual cues into benign prompts to exploit MLLMs’ text-focused safety alignment.
@@
To better improve the attack transferability and its effectiveness, we propose a novel prompt-based jailbreak attack, namely $\mathsf{FC\text{-}Attack}$. Concretely, $\mathsf{FC\text{-}Attack}$ converts harmful queries into harmful flowcharts (images and videos) as visual prompts, allowing users to input benign textual prompts to bypass the model’s safeguards. Specifically, $\mathsf{FC\text{-}Attack}$ consists of two stages: (1) Step-Description Generator Building: In this stage, the step description dataset is synthesized using GPT-4o, and a pre-trained LLM is fine-tuned to obtain a step-description generator. (2) Jailbreak Deployment: This stage uses the generator to produce steps corresponding to the harmful query and generates three types of harmful flowcharts (vertical, horizontal, and S-shaped) as visual prompts. Together with the benign textual prompt, the visual prompt is fed into MLLMs to achieve the jailbreak. Note that the harmful flowcharts are generated automatically without hand-crafted effort.
@@
Our evaluation on the Advbench dataset Zou et al. (2023) shows that $\mathsf{FC\text{-}Attack}$ outperforms previous attacks and achieves an attack success rate (ASR) of over $90\%$ on multiple open-source models, including Llava-Next, Qwen2-VL, and InternVL-2.5, and reaches $94\%$ on the production model Gemini-1.5. Although the ASR is lower on GPT-4o mini, GPT-4o, and Claude-3.5, we how later that it can be improved in certain ways. To further investigate the impact of different elements in flowcharts on the jailbreak effectiveness of MLLMs, we conduct several ablation experiments, including different types of user queries (as shown in Figure˜1), numbers of descriptions, and font styles in flowcharts. These experiments show that MLLMs exhibit higher safety in the text modality but weaker in the visual modality. Moreover, we find that even flowcharts with a one-step harmful description can achieve high ASR, as evidenced by the Gemini-1.5 model, where the ASR reaches $86\%$. Furthermore, font styles in flowcharts also contribute to the ASR increase. For instance, when the font style is changed from “Times New Roman” to “Pacifico”, the ASR increases from $4\%$ to $28\%$ on the model with the lowest ASR (Claude-3.5) under the original style. To mitigate the attack, we consider several popular defense approaches, including Llama-Guard-3-11B-Vision Meta LLaMA (2025), JailGuard Zhang et al. (2024b), AdaShield-S Wang et al. (2024b), and AdaShield-A Wang et al. (2024b). Among them, AdaShield-A demonstrates the best defense performance by reducing the average ASR from $58.6\%$ to $1.7\%$. However, it also reduces MLLM’s utility on benign datasets, which calls for more effective defenses.
@@
In recent years, with the increase in model parameters and training data, LLMs have demonstrated powerful language generation and understanding capabilities Zhao et al. (2023a); Chang et al. (2024), which have driven the emergence of MLLMs Zhang et al. (2024a) (also known as Large Vision Language Models, LVLMs). MLLMs combine visual understanding with language comprehension, showing promising capabilities not only in canonical visual tasks such as Visual Question Answering (VQA) Antol et al. (2015); Khan et al. (2023); Shao et al. (2023), image captioning Hu et al. (2021); Li et al. (2024a), and visual commonsense reasoning Zellers et al. (2019); Tanaka et al. (2021), but also in emerging applications including edited image detection Sun et al. (2025b) and real-time assistance for individuals with visual impairments Zhang et al. (2025). Notably, some MLLMs are capable of processing both image and video inputs, enabling broader applications across multimodal scenarios.
In this paper, we consider both popular open-source and production MLLMs. These MLLMs are the most widely used, and all of them have been aligned to ensure safety. Detailed information is introduced in Appendix˜A.
@@
Similar to LLMs, which have been shown to be vulnerable to jailbreak attacks Yi et al. (2024), MLLMs also remain susceptible despite safety alignment. Current attacks can be categorized into two types: optimization-based and prompt-based attacks. Most existing optimization-based attacks rely on backpropagating the gradient of the target to generate harmful outputs. These methods typically require white-box access to the model, where they obtain the output logits of MLLMs and then compute the loss with the target response to create adversarial perturbations into the visual prompts or textual prompts Bagdasaryan et al. (2023); Shayegani et al. (2024); Qi et al. (2024) (e.g., the target can be “Sure! I’m ready to answer your question.”). Carlini et al. (2023) are the first to propose optimizing input images by using fixed toxic outputs as targets, thereby forcing the model to produce harmful outputs. Building on this, Bailey et al. (2024) introduce the Behaviour Matching Algorithm, which trains adversarial images to make MLLMs output behavior that matches a target in specific contextual inputs. This process requires the model’s output logits to align closely with those of the target behavior. Additionally, they propose Prompt Matching, where images are used to induce the model to respond to specific prompts. Li et al. (2024b) take this further by replacing harmful keywords in the original textual inputs with objects or actions in the image, allowing harmful information to be conveyed through images to achieve jailbreaking. Unlike previous work, these images are generated using diffusion models and are iteratively optimized with models like GPT-4 Achiam et al. (2023). This approach enhances the harmfulness of the images, enabling more effective attacks.
@@
Unlike optimization-based attacks, prompt-based attacks only need black-box access to successfully attack the model without introducing adversarial perturbations into images. Gong et al. (2025) discovers that introducing visual modules may cause the original security mechanisms of LLMs to fail in covering newly added visual content, resulting in potential security vulnerabilities. To address this, they propose the FigStep attack, which converts harmful textual instructions into text embedded in images and uses a neutral textual prompt to guide the model into generating harmful content. This method can effectively attack MLLMs without requiring any training. Wang et al. (2025) identifies a phenomenon named Shuffle Inconsistency, which highlights the tension between “understanding capabilities” and “safety mechanisms” of LLMs. Specifically, even if harmful instructions in text or images are rearranged, MLLMs can still correctly interpret their meaning. However, the safety mechanisms of MLLMs are often more easily bypassed by shuffled harmful inputs than by unshuffled ones, leading to dangerous outputs. Compared to optimization-based attacks, prompt-based attacks usually achieve higher success rates against closed-source models. Our proposed $\mathsf{FC\text{-}Attack}$ also belongs to this category, requiring only black-box access.
@@
Adversary’s Goal. The adversary’s goal is to exploit attacks to bypass the protective mechanisms of MLLMs and access content prohibited by safety policies, e.g., OpenAI’s usage policy OpenAI (2025). This goal takes real-world scenarios into account, where adversaries manipulate the capabilities of MLLMs to easily acquire harmful knowledge and thereby commit criminal acts with minimal learning effort. These objectives pose severe societal impacts and risks to the model providers.
@@
In this section, we introduce the framework of $\mathsf{FC\text{-}Attack}$ (as shown in Figure˜2), which consists of two stages: Step-Description Generator Building and Jailbreak Deployment.
@@
After obtaining the Step-Description Generator $\mathcal{G}$, a harmful query $q_{h}$ is input to generate the corresponding step-by-step description. This description is then processed by a transformation function $\mathcal{F}$ to generate the flowchart (using Graphviz Graphviz Team (2025)). Together with a benign textual prompt $p_{b}$ (more details are in Appendix˜B), the flowchart will be fed into the aligned MLLM $\mathcal{A}$ to produce the harmful output $o_{h}$, as shown below:
@@
Target Model. We test $\mathsf{FC\text{-}Attack}$ on seven popular MLLMs, including the open-source models Llava-Next (llama3-llava-next-8b) Liu et al. (2024c), Qwen2-VL (Qwen2-VL-7B-Instruct) Wang et al. (2024a), and InternVL-2.5 (InternVL-2.5-8B) Chen et al. (2024a) as well as the production models GPT-4o mini (gpt-4o-mini-2024-07-18) OpenAI (2024), GPT-4o (gpt-4o-2024-08-06) Hurst et al. (2024), Claude (claude-3-5-sonnet-20240620) Anthropic (2024), and Gemini (gemini-1.5-flash) Google (2024). Moreover, we also test $\mathsf{FC\text{-}Attack}$ on MLLMs via video, including Qwen-VL-Max (Qwen-VL-Max-latest) Bai et al. (2025b), Qwen2.5-Omni Xu et al. (2025) and LLaVA-Video-7B-Qwen2 Zhang et al. (2024d).
Dataset. Following Chao et al. (2025b), we utilize the deduplicated version of Advbench Zou et al. (2023), which includes 50 representative harmful queries. Based on Advbench, we use $\mathsf{FC\text{-}Attack}$ to generate $3$ types of flowcharts for each harmful query, which includes $150$ jailbreak flowcharts in total. To assess whether defense methods have the critical issue of “over-defensiveness” when applied to benign datasets, we utilize a popular evaluation benchmark, MM-Vet Yu et al. (2024).
@@
HADES 44 $16$ 22 22 2020 $10$ 88
SI-Attack $36$ $14$ 0 $69$ $24$ $42$ $40$
$M$-SafetyBench 0 0 0 $50$ $50$ $54$ $16$
VA-Jailbreak 66 $18$ 22 22 $40$ $22$ $16$
FigStep 0 22 0 $30$ $62$ $36$ 0
Ours (Vertical) 88 88 0 $76$ $76$ $84$ $68$
@@
$\text{ASR}=\frac{\text{\# Queries Successfully Jailbroken}}{\text{\# Original Harmful Queries}}.$ (5)
Following the judge prompt Chao et al. (2025b), we employ GPT-4o to serve as the evaluator.
FC-Attack Deployment. Referring to Section˜4, $\mathsf{FC\text{-}Attack}$ consists of two stages. For the Step-Description Generator Building, we first use GPT-4o to randomly generate several daily topics and $3$ few-shot examples, which are then combined into a prompt and fed into GPT-4o to construct the dataset $\mathcal{D}_{t}$. In our experiments, the number of descriptions in the flowchart is limited to a maximum of $10$ steps, as too many descriptions can result in excessive length in one direction of the image. The dataset contains $5,000$ pairs of queries and step descriptions for daily activities, with the temperature set to 1 (more details are provided in Appendix˜C). We then select Mistral-7B-Instruct-v0.1 Jiang et al. (2023) as the pre-trained LLM and fine-tune it on $\mathcal{D}_{t}$ using LoRA. The fine-tuning parameters include a rank of $16$, a LoRA alpha value of $64$, 22 epochs, a batch size of $8$, a learning rate of 1e51e-5, and a weight decay of 1e51e-5. For the jailbreak deployment stage, we set the temperature to $0.3$ for all MLLMs for a fair comparison.
Baselines. To validate the effectiveness of $\mathsf{FC\text{-}Attack}$, we adopt five jailbreak attacks as baselines, which are categorized into black-box attacks (MM-SafetyBench Liu et al. (2024d), SI-Attack Zhao et al. (2025), and FigStep Gong et al. (2025)) and white-box attacks (HADES Li et al. (2024b), VA-Jailbreak Qi et al. (2024)).
For black-box attacks, MM-SafetyBench utilizes StableDiffusion Rombach et al. (2022) and GPT-4 Achiam et al. (2023) to generate harmful images and texts based on Advbench. The input harmful images and texts used in SI-Attack are from the outputs of MM-SafetyBench, while FigStep is set up using their default settings Gong et al. (2025).
For white-box attacks, all input data, including images and texts, is obtained from MM-SafetyBench’s outputs, with the attack step size uniformly set to 1/255. HADES employs LLaVa-1.5-7b Liu et al. (2024b) as the attack model, running 3,000 optimization iterations with a batch size of 2. For VA-Jailbreak, LLaVa-1.5-7b Liu et al. (2024b) is used as the attack model, setting the epsilon of the attack budget to 32/255, with 5,000 optimization iterations and a batch size of 8. To align with the black-box scenario considered in this paper, we adopt a model transfer strategy, where these white-box methods are trained on one model (LLava-1.5-7b) and then transferred to our target testing models.
@@
To mitigate the attacks, we explore several possible defense methods including Llama-Guard3-V, JailGuard, ECSO and AdaShield. Llama-Guard3-V (Llama-Guard-3-11B-Vision) Meta LLaMA (2025) determines whether the input is safe by feeding both the image and text into the model. JailGuard Zhang et al. (2024b) generates input variants and evaluates them using MiniGPT-4 Zhu et al. (2024), identifying harmful content by comparing differences in the responses. ECSO Chao et al. (2025a) is less intrusive but mainains a high safety rate simultaneously. AdaShield-S employs static prompts in the textual prompt to defend against attacks, while AdaShield-A uses Vicunav1.5-13B as a defender to adaptively rewrite defensive prompts Wang et al. (2024b).
@@
In this section, we explore the performance of $\mathsf{FC\text{-}Attack}$ and conduct ablation study and defense research. We conduct jailbreak experiments on MLLMs for $\mathsf{FC\text{-}Attack}$. As shown in Figure˜3 , it is a successful jailbreak case on Gemini-1.5.
Figure 3: Successful jailbreak instance on Gemini-1.5 using $\mathsf{FC\text{-}Attack}$.
@@
[…]
@@
Jailbreaking via Images. In Table˜1, we compare the performance of $\mathsf{FC\text{-}Attack}$ with different baseline methods on both open-source and production models. We observe that $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves the highest ASR on both models compared to all baselines. For example, the ASRs are $94\%$, $92\%$, $90\%$, and $90\%$ on Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, respectively. However, the ASR on some production models, such as Claude-3.5, GPT-4o, and GPT-4o mini, is relatively low, at $4\%$, $30\%$, and $10\%$, respectively. This might be because these production models have more advanced and updated visual safety alignment strategies.
@@
Jailbreaking via Videos. To conduct attacks from the video modality, we transform each jailbreak image into a 3-second video by setting all frames into the same image. Note that we also consider the Procedure Flowcharts, where each part (1 question and 5 steps) has been sequentially filled into a 0.5s video frame, resulting in a 3s video. We then evaluate the effectiveness of video jailbreak on three models: Qwen-VL-Max, Qwen2.5-Omni and LLaVA-Video. The performance is summarized in Table˜4. Our $\mathsf{FC\text{-}Attack}$ (Ensemble) achieves a stable 88% ASR, whereas HADES peaks at 46% on Qwen-VL-Max (dropping to 28% on LLaVA-Video) and Figstep fluctuates between 78% on Qwen-VL-Max and 2% on Qwen2.5-Omni, highlighting our method’s consistent performance across models. As shown in Figure˜A2, attacks using harmful text have an extremely low ASR. When the same harmful queries and steps are delivered via the video modality, the MLLMs become highly vulnerable, with ASR up to 88%.
@@
Figure 4: ASR under different prompts against MLLMs.
As shown in Figure˜4 when using only the harmful query (text) as input, we observe very low ASR. The ASR is $0\%$ on GPT-4o mini, GPT-4o, Claude-3.5, Qwen2-VL, and InternVL-2.5, and only $2\%$ and $6\%$ on Gemini-1.5 and Llava-Next, respectively. This indicates that the textual modality of these MLLMs has relatively robust defenses against such inputs. However, when both the harmful query and the step descriptions are input as text, the ASR increases to $36\%$ on Gemini-1.5, and to $16\%$ and $6\%$ on Llava-Next and InternVL-2.5, respectively, while remaining at $0\%$ on the other models. When this information is converted into a flowchart and only a benign textual prompt is provided, the ASR on these models improves significantly. This demonstrates that the defenses of MLLMs in the visual modality have noticeable weaknesses compared with the language modality.
@@
Numbers of Steps in Flowcharts. As described in Section˜4, flowcharts of $\mathsf{FC\text{-}Attack}$ are generated from step descriptions. In this section, we aim to explore the impact of the number of steps in flowcharts on jailbreak effectiveness. Therefore, we reduce the number of steps to 1, 3, and 5, respectively. Table˜2 presents the ASR results for four types of flowcharts (Vertical, Horizontal, S-shaped, and Ensemble) with varying numbers of steps. We find that, even with only one step in the description, flowcharts achieve relatively high ASR. For example, for Gemini-1.5, Llava-Next, Qwen2-VL, and InternVL-2.5, the ASR for Ensemble at 1 step is $86\%$, $70\%$, $88\%$, and $82\%$, respectively. As the number of steps increases, the ASR for almost all flowchart types improves significantly. For instance, the Horizontal ASR of Gemini-1.5 increases from $78\%$ at “1 step” to $86\%$ at “3 steps” and $88\%$ at “5 steps”. Similarly, the S-shaped ASR of InternVL-2.5 improves from $68\%$ at “1 step” to $92\%$ at “5 steps”. This suggests that increasing the number of step descriptions makes the model more vulnerable and susceptible to jailbreak attacks.